# https://www.outerlands.io/

### v2

@@ 3 Search costs
There are values of $s_{A}$, $s_{V}$, $z$, and $x$ that admit a mixed equilibrium that is not the same as that found in Lemma 2.1 if both $A$ and $V$ are rational and risk neutral.
@@
We see that this constraint on be conforms to the requirement that $10<b<$ 1. Thus there are values for $g$, $h$, and $b$ all strictly between 00 and $1$ that admit an equilibrium. This is not the pure strategy equilibrium from Lemma 2.1 that is intended by the optimistic design. ∎
@@ 3.1 The Role of the Transactor
What we have shown is that strategies other than those we found in Lemma 2.1 constitute the equilibrium. We have found constraints on the model values such that if those constraints are met, the validator does no search $b$ of the time and, given no search, blindly challenges $g$ of the time. The validator is honest only $h$ of the time. The transactor will still participate despite this added uncertainty. The precise values of these depend on the values of the search cost, the stakes, and the potential benefit of getting away with dishonest behavior.
@@ 3.2 Remark
4 Potential Solutions
The verifier’s dilemma has yet to become a practical problem because existing optimistic rollups have yet to decentralize their sequencers. Keeping this status quo is one potential solution, but asking how decentralization might be achieved without compromising the security of the chain is an important question. Another trivial solution would be to revert to a mechanism that requires the validators to check the integrity of the block every time, as is done with proof of stake L1s and zero knowledge (ZK) rollups. As ZK rollups get faster, this may become a more attractive option. We present here the outline of potential solutions that are more in keeping with the spirit of optimistic rollups.
4.1 Random Checks
If we were to stipulate that each block be checked, we would stray away from the spirit of optimistic rollups and lose their lightweight efficiency. Requiring intermittent, random checks could keep the optimistic efficiency without the hazard of the verifier’s dilemma. Whether such a set up would indeed be meaningfully more efficient than simply checking every time would depend on the relative value of the potential benefit from attack and the potential penalties. It would also depend on precisely how the random checks would be implemented. This would not necessitate fraud proofs, which are needed to find precisely where the chain became faulty, and instead would require proofs of just the block in question. These proofs could be done by nodes, which would again need incentives, or by smart contracts on the L1, which could have expensive execution that might fall on the users or be paid for with inflation. If we assume this could be done with a smart contract without too much additional cost to the user, we get a game where $\mathcal{C}$ is the contract ‘player’ in the place of $V$. Let us also define $p$, where $10<p<$ 1, to be the probability with which the contract checks the proof.
Lemma 5
There are values of the probability of check, $p$, $s_{A}$ and $z$ such that a rational, risk neutral aggregator will always be honest.
{istgame}
[…]
Game 3: Game tree illustrating a random check where utility is that of $A$ and the dotted line indicates an information set.
It is immediate that the expected income on the left child of the root is zero. Recall this is because we normalized it so. Since the left, honest choice on the part of $A$ is the desired outcome, we need a $p$ such that the right side yields negative income in expectation. Thus we would need,
$(1-p)z-ps_{A}<0$
$(1-p)z<ps_{A}$
Solving for $p$ yields,
$ps_{A}>z-pz$
$p(s_{A}+z)>z$
$p>\frac{z}{z+s_{A}}$
This is clearly in keeping with our requirement that $10<p<$ 1, but is also clear that, given $z>s_{A}$ and especially in cases where $z$ is much larger than $s_{A}$, $p$ will be close to $1$. Thus, there are only minor savings as compared to checking every time, but there are values for which such a setup would result in a rational, risk neutral $A$ behaving honestly, as desired. ∎
In this proof, we have assumed that there are only random checkers and no traditional validators. This is a viable solution, although it requires the approximate value of $z$ to be known and for the smart contract to work as described. A scenario with both random and traditional nodes might yield a different, potentially more efficient, result.
4.2 Random Mistakes
As we have shown in the proof for Eq. 2, there is an expression for $h$ for which the validator is indifferent between search and no search. Thus, we can presume that there are still values of $b$ and $g$ between 00 and $1$ that will yield a mixed equilibrium different from that desired. Modeling this precisely would require coefficients for both whether the block is being produced by a mistake-making $A$ and, if so, whether it elects to make a mistake. A non-mistake-making $A$ would have to take these terms into account. We leave the realization of such a model to future work, but it is far from immediate that reasonable values would yield an equilibrium in which the validator searches every time. It would likely be unreasonable for a mistake-making $A$ to make mistakes every time, as this would be tremendously expensive.
4.3 Easter Eggs
If the search costs, $x$, are relatively small, compensating the validators directly for their services seems like a tempting solution. However, it may not be clear whether or not the search is being performed correctly and checking this search would be yet another role with its own incentives. If it were possible to hide reward ‘Easter eggs,’ which could be found through proper searching but not a random challenge, this could act as a delivery mechanism for validator rewards. These rewards would be realized even when the aggregator behaves honestly.
Lemma 6
(Sketch) Compensating the validator with some term $e\geq x$, will result in the expected income for the validator of always searching being higher than not searching.
Let us add $e$, the expected value of finding the Easter egg, to the validator income in the case of search. Reading off from the proof of Eq. 2, we have that search yields more income if
$(1-h)\frac{1}{2}s_{A}-x+e>-ghs_{V}+g(1-h)\frac{1}{2}$
since $g<1$, it is clear to see that if we have $-x+e>-ghs_{V}$, the above term is always true. If we require that $e\geq x$, the result is immediate. ∎

### v3

@@ 4.3 Easter Eggs
There are values of $s_{A}$, $s_{V}$, $z$, and $x$ that admit a mixed equilibrium that is not the same as that found in Lemma 2 if both $A$ and $V$ are rational and risk neutral.
@@
We see that this constraint on be conforms to the requirement that $10<b<$ 1. Thus there are values for $g$, $h$, and $b$ all strictly between 00 and $1$ that admit an equilibrium. This is not the pure strategy equilibrium from Lemma 2 that is intended by the optimistic design. ∎
@@ 4.4 The Role of the Transactor
What we have shown is that strategies other than those we found in Lemma 2 constitute the equilibrium. We have found constraints on the model values such that if those constraints are met, the validator does no search $b$ of the time and, given no search, blindly challenges $g$ of the time. The validator is honest only $h$ of the time. The transactor will still participate despite this added uncertainty. The precise values of these depend on the values of the search cost, the stakes, and the potential benefit of getting away with dishonest behavior.